how many grams of AgNO3 needed to make 500mL of 0.25 M solution?

Grade:10

3 Answers

Shreyash Rane
7 years ago
21250 grams.1st apply formula M=no. Of moles/volume.of soln.AgNO3=170grams and so eventually find the answer21250 grams
Shreyash Rane
7 years ago
21250 grams.1st apply formula M=no. Of moles/volume.of soln.AgNO3=170grams and so eventually find the answer21250 gramsBest of luck
Goutham
6 years ago
The solution should contain:
0.125 moles of AgNO3  (0.25 moles in one litre => 0.125 moles in 500 mL)
number of moles*Atomic weight= Given weight
0.125*(10+14+(16*3))=0.125*170 = 21.250g